### v1

# Fractal counterparts of density combinatorics theorems in continued fractions

Abstract.
We build a bridge from density combinatorics to dimension theory of continued fractions. We establish a fractal transference principle that transfers common properties of subsets of $\mathbb{N}$ with positive upper density to properties of subsets of irrationals in $(0,1)$ for which the set $\{a_{n}(x)\colon n\in\mathbb{N}\}$ of partial quotients induces an injection $n\in\mathbb{N}\mapsto a_{n}(x)\in\mathbb{N}$. Let ()(*)( ∗ ) be a certain property that holds for any subset of $\mathbb{N}$ with positive upper density. The principle asserts that for any subset $S$ of $\mathbb{N}$ with positive upper density, there exists a set $E_{S}$ of Hausdorff dimension $1/2$ such that the set $\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S}}\{a_{n}(x)\}\cap S$ has the same upper density as that of $S$, and thus inherits property ()(*)( ∗ ). Examples of ()(*)( ∗ ) include the existence of arithmetic progressions of arbitrary lengths and the existence of arbitrary polynomial progressions, known as Szemerédi’s and Bergelson-Leibman’s theorems respectively. In the same spirit, we establish a relativized version of the principle applicable to the primes, to the primes of the form $y^{2}+z^{2}+1$, to the sets given by the Piatetski-Shapiro sequences.
2020 Mathematics Subject Classification:
05D10, 11A55, 11K55
Keywords: arithmetic progression; continued fraction; Hausdorff dimension
1. Introduction
Problems on the existence of arithmetic progressions lie at the interface of a number of fields, including combinatorics, number theory, ergodic theory and dynamical systems. Van der Waerden’s theorem [39] states that if the set $\mathbb{N}$ of positive integers is partitioned into finitely many sets, then one of these sets contains an $\ell$-term arithmetic progression for every $\ell\geq 3$. A significant strengthening of Van der Waerden’s theorem was obtained by Szemerédi [35] who proved that any subset of $\mathbb{N}$ with positive upper density contains an $\ell$-term arithmetic progression for every $\ell\geq 3$. For a set $S\subset\mathbb{N}$, define its upper density by
$\overline{d}(S)=\limsup_{N\to\infty}\frac{1}{N}\#({S\cap[1,N]}),$
and its upper Banach density by
$d_{\rm B}(S)=\limsup_{N\to\infty}\frac{1}{N}\sup_{M\in\mathbb{N}}\#({S\cap[M,M+N]}).$
Note that $d_{\rm B}(S)\geq\overline{d}(S)$. The positive upper density assumption in Szemerédi’s theorem can actually be weakened to the positivity of upper Banach density.
Furstenberg [10, 11] established the multiple recurrence theorem and the correspondence principle, linking the field of density combinatorics to recurrence in ergodic theory. Using this principle he obtained a new proof of Szemerédi’s theorem. This development laid a foundation for further extensions of Szemeredi’s theorem to several different directions, see [2, 12, 14, 34, 36] for example. The aim of this paper is to provide counterparts of these density combinatorics theorems in dimension theory of continued fractions.
Each $x\in(0,1)\setminus\mathbb{Q}$ has the unique infinite regular continued fraction expansion
[…]
where each positive integer $a_{n}(x)$, $n\in\mathbb{N}$ is called a partial quotient or simply a digit of $x$. We analyze the structure of the set of digits as an infinite subset of $\mathbb{N}$. For this purpose, overlapping digits do not count and we may restrict to the set of points whose digits are all different:
$E=\{x\in(0,1)\setminus\mathbb{Q}\colon a_{m}(x)\neq a_{n}(x)\ \text{\rm for all }m,n\in\mathbb{N},\ m\neq n\}.$
An investigation of this set was suggested by Erdős in the early stage of dimension theory of continued fractions. Ramharter [31] proved that $E$ is of Hausdorff dimension $1/2$. Ramharter’s result is a refinement of Good’s landmark theorem [13] which states that the set $\{x\in(0,1)\setminus\mathbb{Q}\colon a_{n}(x)\to\infty\text{ as }n\to\infty\}$ is of Hausdorff dimension $1/2$.
1.1. Statements of results and their consequences
We now state our two main results (Theorem 1.1 and Theorem 1.4), along with five consequential theorems which are manifestations of density combinatorics theorems in the regular continued fraction.
1.1.1. First main result
Let $\dim_{\rm H}$ denote the Hausdorff dimension on $[0,1]$.
Theorem 1.1 (fractal transference principle).
Let $S\subset\mathbb{N}$.
  • (a)
    If $\overline{d}(S)>0$, then there exists $E_{S}\subset E$ such that
    $\dim_{\rm H}E_{S}=\dim_{\rm H}E=\frac{1}{2}\ \text{ and }\ \overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S}}\{a_{n}(x)\}\cap S\right)=\overline{d}(S).$ (1.2)
    In particular,
    $\dim_{\rm H}\left\{x\in E\colon\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)\right\}=\dim_{\rm H}E.$ (1.3)
  • (b)
    If $d_{\rm B}(S)>0$, then there exists $F_{S}\subset E$ such that
    $\dim_{\rm H}F_{S}=\dim_{\rm H}E\ \text{ and }\ d_{\rm B}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\right)=d_{\rm B}(S).$ (1.4)
    In particular,
    $\dim_{\rm H}\left\{x\in E\colon d_{\rm B}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=d_{\rm B}(S)\right\}=\dim_{\rm H}E.$ (1.5)
Equation (1.3) is a ‘pointwise version’ of the second equation in (1.2). Equation (1.3) asserts that the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)$ does not cause a dimension drop in $E$. More importantly, it allows us to immediately transfer statements on subsets of $\mathbb{N}$ with positive upper density to statements on continued fractions of points in sets of Hausdorff dimension $1/2$. For example, from (1.3) and Szemerédi’s theorem it follows that if $S\subset\mathbb{N}$ has positive upper density then for any $x\in(0,1)\setminus\mathbb{Q}$ in a set of Hausdorff dimension $1/2$ and for every integer $\ell\geq 3$, there exist $k(x)\in\mathbb{N}$, $m(x)\in\mathbb{N}$, $n_{1}(x),n_{2}(x),\ldots,n_{\ell}(x)\in\mathbb{N}$ such that $k(x)+m(x),k(x)+2m(x),\ldots,k(x)+\ell m(x)\in S$ and $a_{n_{1}}(x)=k(x)+m(x),a_{n_{2}}(x)=k(x)+2m(x),\ldots,a_{n_{\ell}}(x)=k(x)+\ell m(x)$. Equations (1.2) ensure that all these integers are taken uniformly on the set $E_{S}$ of Hausdorff dimension $1/2$. Moreover, from the uniformity in multiple recurrence (see [1, Theorem F2]) and Furstenberg’s correspondence principle, the common difference $m(x)$ $(x\in E_{S})$ of the $\ell$-term arithmetic progression can be taken to be bounded from above by a universal constant that depends only on $\ell$ and $\overline{d}(S)$. Even if the upper density of $S$ is zero, similar statements follow from (1.4) and (1.5) provided $S$ has positive upper Banach density.
1.1.2. Polynomial progressions in $\mathbb{N}$
We proceed to giving more examples of uniform transfer via Theorem 1.1. Bergelson and Leibman established a multiple recurrence theorem with polynomial times [2, Theorem ${\rm A}_{0}$], and as its corollary derived a density combinatorics theorem [2, Theorem ${\rm B}_{0}$] which in particular asserts that any subset of $\mathbb{Z}$ with positive upper Banach density contains ‘arbitrary polynomial progressions’. Their theorem is a vast generalization of Szemerédi’s theorem. For an application to the continued fraction, we need the following version of Bergelson-Leibman’s theorem in $\mathbb{N}$. For completeness we include a proof in Appendix A. Let $\mathbb{Z}_{0}[X]$ denote the set of polynomials with integer coefficients that vanish at $X=0$.
Theorem 1.2 (cf. [2] Theorem ${\rm B}_{0}$).
If $S\subset\mathbb{N}$ and $d_{\rm B}(S)>0$, then for any finite collection $h_{1},h_{2}\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{N}$, $m\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S.$
Transferring Theorem 1.2 to the continued fraction via Theorem 1.1 we obtain the following statement.
Theorem 1.3.
If $S\subset\mathbb{N}$ and $\overline{d}(S)>0$ (resp. $d_{\rm B}(S)>0$), then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist [^1] A close inspection into the general construction in §$\S$3.2 shows that in the $\ell$-term arithmetic progression case $h_{1}(X)=X,h_{2}(X)=2X,\ldots,h_{\ell}(X)=\ell X$ one can choose the indices $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ so that $n_{1}<n_{2}<\cdots<n_{\ell}$. Similar remarks apply to Theorems 1.6, 1.9 and 1.11. $k\in\mathbb{N}$, $m\in\mathbb{N}$, $n_{1},n_{2}\ldots,n_{\ell}\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S$ and $a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m)$ for any $x\in E_{S}$ (resp. $x\in F_{S}$).
1.1.3. Second main result
Theorem 1.3 does not apply to subsets of $\mathbb{N}$ whose upper Banach density is zero. To accommodate such sets, notably the set of primes, we develop a relativized version of Theorem 1.1. Given two infinite sets $A\subset S\subset\mathbb{N}$, the upper density of $A$ relative to $S$ is defined by
$\overline{d}(A|S)=\limsup_{N\to\infty}\frac{\#(A\cap[1,N])}{\#(S\cap[1,N])}.$
We say $S$ has polynomial density if there exist constants $C_{1}>0$, $C_{2}>0$ and $\alpha\geq 1$, $\beta\geq 0$ such that for all sufficiently large $N>1$,
$\frac{C_{1}N^{1/\alpha}}{(\log N)^{\beta}}\leq\#(S\cap[1,N])\leq\frac{C_{2}N^{1/\alpha}}{(\log N)^{\beta}}.$
In this case we also say $S$ has polynomial density with exponent $\alpha$. Our second main result is a relativized version of Theorem 1.1.
Theorem 1.4 (relative fractal transference principle).
Let $S$ be an infinite subset of $\mathbb{N}$ that has polynomial density with exponent $\alpha$. For any $A\subset S$ with $\overline{d}(A|S)>0$, there exists a subset $E_{S,A}$ of $\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$ such that
$\dim_{\rm H}E_{S,A}=\dim_{\rm H}\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}=\frac{1}{2\alpha}$ (1.6)
and
$\overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A|S\right)=\overline{d}(A|S).$ (1.7)
In particular,
[…]
dimH{ :xE:xE (1.8)
Since $\mathbb{N}$ has polynomial density with $\alpha=1$, $\beta=0$, Theorem 1.4 includes Theorem 1.1(a) as a special case. Equation (1.8) is a ‘pointwise version’ of the first equality in (1.6), and allows us to immediately transfer density combinatorics statements on subsets of $S$ to statements in dimension theory of continued fractions. This transfer can be done uniformly on $E_{S,A}$ by virtue of (1.7), as developed below.
1.1.4. Polynomial progressions in primes
The first application of Theorem 1.4 is to the set $\mathbb{P}$ of prime numbers. It is well-known (see e.g., [23, Corollary 3.4]) that the Banach density of $\mathbb{P}$ is zero. Green and Tao [14, Theorem 1.2] proved that any subset of $\mathbb{P}$ with positive upper relative density contains an $\ell$-term arithmetic progression for every $\ell\geq 3$, which they call ‘Szemerédi’s theorem in prime numbers’. Green-Tao’s theorem is a corollary of the wide open conjecture of Erdős and Turán [7] which states that any subset of $\mathbb{N}$ with divergent reciprocal sum contains arithmetic progressions of arbitrary lengths. Subsequently, Tao and Ziegler [36] extended Green-Tao’s theorem to polynomial progressions as follows.
Theorem 1.5 ([36] Theorem 1.3).
If $A\subset\mathbb{P}$ and $\overline{d}(A|\mathbb{P})>0$, then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in A.$
From the prime number theorem it follows that $\mathbb{P}$ has polynomial density with $\alpha=\beta=1$. Transferring Theorem 1.5 to the continued fraction via Theorem 1.4 we obtain the following statement.
Theorem 1.6.
If $A\subset\mathbb{P}$ and $\overline{d}(A|\mathbb{P})>0$, then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in A$ and $a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m)$ for any $x\in E_{\mathbb{P},A}.$
Taking $A=\mathbb{P}$ in Theorem  1.6 yields the following statement.
Corollary 1.7.
We have
[…]
Surprisingly, there are few literature on continued fractions with prime digits. Schindler and Zweimüller [33] investigated probabilistic laws governing the appearance of prime regular continued fraction digits for irrationals which are typical in the sense of the Lebesgue measure on $(0,1)$. Das and Simmons [4] investigated the set of irrationals in $(0,1)$ all whose regular continued fraction digits are prime numbers.
1.1.5. Arithmetic progressions in primes of quadratic form
We give three more examples of application of Theorem 1.4, one of which is still related to prime numbers. Let $\mathbb{P}(1)$ denote the set of prime numbers represented by a quadratic form $x^{2}+y^{2}+1$, i.e.,
$\mathbb{P}(1)=\{p\in\mathbb{P}\colon p=x^{2}+y^{2}+1\ \text{\rm for some }x,y\in\mathbb{N}\}.$
Linnik [20] proved that $\mathbb{P}(1)$ contains infinitely many primes. Motohashi [25] proved an upper bound #((1)[$\#(\mathbb{P}(1)\cap[1,N])\leq CN/(\log N)^{\frac{3}{2}}$, and Iwaniec [17, Theorem 1] obtained a lower bound of the same order, namely $\mathbb{P}(1)$ has polynomial density with $\alpha=1$, $\beta=3/2$. For more information on $\mathbb{P}(1)$, see [21, 41] for example.
Teräväinen [37] proved that ‘almost every’ even positive integer $n\not\equiv 5,8$ mod $9$ can be written as the sum of two elements of $\mathbb{P}(1)$. As a by-product of a proof of this, he also proved that any subset of $\{p\in\mathbb{P}\colon p=x^{2}+y^{2}+1\ \text{\rm for some coprime }x,y\in\mathbb{N}\}$ with positive relative upper density contains infinitely many non-trivial $3$-term arithmetic progressions. Sun and Pan [34] extended Green-Tao’s theorem to $\mathbb{P}(1)$.
Theorem 1.8 ([34] Theorem 1.1).
If $A\subset\mathbb{P}(1)$ and $\overline{d}(A|\mathbb{P}(1))>0$, then $A$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.
Transferring Theorem 1.8 to the continued fraction via Theorem 1.4 we obtain the following statement.
Theorem 1.9.
If $A\subset\mathbb{P}(1)$ and $\overline{d}(A|\mathbb{P}(1))>0$, then for any integer $\ell\geq 3$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+m,k+2m,\ldots,k+\ell m\in A$ and $a_{n_{1}}(x)=k+m,a_{n_{2}}(x)=k+2m,\ldots,a_{n_{\ell}}(x)=k+\ell m$ for any $x\in E_{\mathbb{P}(1),A}.$
1.1.6. Arithmetic progressions in sets given by Piatetski-Shapiro sequences
For $x\geq 0$ let $\lfloor x\rfloor$ denote the largest integer not exceeding $x$. For $\alpha>1$ we consider the set
${\rm PS}(\alpha)=\{\lfloor n^{\alpha}\rfloor\colon n\in\mathbb{N}\}.$
For $\alpha>1$ and $\alpha\notin\mathbb{N}$, the sequence $(\lfloor n^{\alpha}\rfloor)_{n=1}^{\infty}$ is called Piatetski-Shapiro’s sequence [29], after Piatetski-Shapiro [29] who proved that for all $111<\alpha<12/$ 11 the sequences contain infinitely many prime numbers. In spite of a vast amount of literature, prime numbers in Piatetski-Shapiro’s sequences are not a topic of this paper. There is no integer $\alpha>1$ such that ${\rm PS}(\alpha)$ contains arithmetic progressions of arbitrary lengths [3, 5, 6]. Since ${\rm PS}(\alpha)$ has a convergent reciprocal sum, it is independent of the conjecture of Erdős and Turán.
Piatetski-Shapiro’s sequence is obtained by curving with bounded errors the sequence $(n^{\alpha})_{n=1}^{\infty}$ whose graph $\{(n,n^{\alpha})\in\mathbb{R}^{2}\colon n\in\mathbb{N}\}$ is contained in the graph of the smooth function $x\in[1,\infty)\mapsto x^{\alpha}$. Positive integer sequences with this kind of properties are called slightly curved sequences [32]. Saito and Yoshida [32, Theorem 4] extended Szemerédi’s theorem to graphs of a class of slightly curved sequences, and as a corollary proved that for every $21<\alpha<$ 2, the graph $\{(n,\lfloor n^{\alpha}\rfloor)\in\mathbb{N}^{2}\colon n\in\mathbb{N}\}$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$ (see [32, Corollary 6]). From this it follows that ${\rm PS}(\alpha)$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$. In fact, the next theorem follows from combining the results in [32]. For completeness we include a proof in Appendix B.
Theorem 1.10.
Let $21<\alpha<$ 2. If $A\subset{\rm PS}(\alpha)$ and $\overline{d}(A|{\rm PS}(\alpha))>0$, then $A$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.
Clearly ${\rm PS}(\alpha)$ has polynomial density with exponent $\alpha$ and $\beta=0$. Transferring Theorem 1.10 to the continued fraction via Theorem 1.4 we obtain the following statement.
Theorem 1.11.
Let $21<\alpha<$ 2. If $A\subset{\rm PS}(\alpha)$ and $\overline{d}(A|{\rm PS}(\alpha))>0$, then for any integer $\ell\geq 3$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+m,k+2m,\ldots,k+\ell m\in A$ and $a_{n_{1}}(x)=k+m,a_{n_{2}}(x)=k+2m\ldots,a_{n_{\ell}}(x)=k+\ell m$ for any $x\in E_{{\rm PS}(\alpha),A}.$
1.1.7. Polynomial progressions in subsets of $\mathbb{N}$ with positive lower density
The lower density of $S\subset\mathbb{N}$ is given by
$\underline{d}(S)=\liminf_{N\to\infty}\frac{1}{N}\#({S\cap[1,N]}).$
Any subset of $\mathbb{N}$ with positive lower density has polynomial density with $\alpha=1$, $\beta=0$. Transferring Theorem 1.2 to the continued fraction via Theorem 1.4 we obtain the following statement.
Theorem 1.12.
Let $S\subset\mathbb{N}$ satisfy $\underline{d}(S)>0$. If $A\subset S$ and $\overline{d}(A|S)>0$, then for any finite collection $h_{1},h_{2}\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S$ and $a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m)$ for any $x\in E_{S,A}$.
1.2. Essential role of non-monotonicity of digits
Arithmetic progressions in the continued fraction was first studied by Tong and Wang [38], who considered the smaller set
$J=\{x\in(0,1)\setminus\mathbb{Q}\colon a_{n}(x)<a_{n+1}(x)\ \text{ for every }n\geq 1\}\subset E,$
and proved that the set of $x\in J$ whose digit set $\{a_{n}(x)\colon n\in\mathbb{N}\}$ contains an $\ell$-term arithmetic progression with a common difference $m$ for every $\ell\geq 3$ and every $m\geq 1$ is of Hausdorff dimension $1/2$. The argument in [38] is constructive and does not rely on Szemerédi’s theorem. Zhang and Cao [42] proved that the set of points in $J$ whose digit sets have upper Banach density $1$ is of Hausdorff dimension $1/2$, a statement comparable to (1.5) in the special case $S=\mathbb{N}$.
Since the map $x\in J\mapsto\{a_{n}(x)\colon n\in\mathbb{N}\}\in 2^{\mathbb{N}}$ induces a bijection between $J$ and the collection of infinite subsets of $\mathbb{N}$, $J$ is indeed a natural set to investigate. However, the next theorem of independent interest, to be proved in §$\S$4.3, suggests that $E$ rather than $J$ is the right choice for our purpose of building a bridge from density combinatorics to dimension theory of continued fractions.
Theorem 1.13.
Let $S\subset\mathbb{N}$.
  • (a)
    If $\overline{d}(S)>0$, then we have
    $\dim_{\rm H}\left\{x\in J\colon\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)\right\}=0<\dim_{\rm H}J.$
  • (b)
    If $S$ has polynomial density with exponent $\alpha<2$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ we have
[…]
    dimH{ :xJ:xJ
In sharp contrast to (1.3), Theorem 1.13(a) asserts that for any $S\subset\mathbb{N}$ with positive upper density, the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)$ forces a significant dimension drop in $J$. In contrast to (1.8), Theorem 1.13(b) asserts that the conditions $\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S$ and $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)=\overline{d}(A|S)$ together force a dimension drop in $J$. The Hausdorff dimension drops to zero in the case $\alpha=1$, such as $S=\mathbb{P}$ and $S=\mathbb{P}(1)$. These two comparisons show that the non-monotonicity of continued fraction digits for points in $E\setminus J$ is crucial in both Theorems 1.1 and 1.4.
1.3. Outline of proofs of the main results
Notice that Theorem 1.1(a) follows from Theorem 1.4 since $\overline{d}(S|\mathbb{N})=\overline{d}(S)$ holds for any $S\subset\mathbb{N}$. Moreover, a proof of Theorem 1.1(b) and that of Theorem 1.4 rely on the same set of ideas. Hence, we only outline the proof of Theorem 1.4 below.
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha\geq 1$, and let $A\subset S$ have positive upper density relative to $S$. The construction of the set $E_{S,A}$ in Theorem 1.4 breaks into two main steps. The central idea is to use Moran-like fractals. A Moran fractal is a Cantor-like set with a tractable structure, and has been used extensively for estimating Hausdorff dimension, see [22, 24, 27, 28] for example. In the first step we choose an infinite subset $S_{*}$ of $S$, and construct a Moran-like fractal of Hausdorff dimension $1/(2\alpha)$ consisting of points whose digit sets are contained in $S_{*}$. This set is our seed set, with no relation to $A$. The second step is to modify the seed set into another Moran-like fractal with exactly the same Hausdorff dimension, by adding elements of $A$ into the digit sequences of points in the seed set. The resultant Moran-like fractal is our target set $E_{S,A}$.
The two-step construction of $E_{S,A}$ has been inspired by the afore-mentioned papers [38, 42] that construct subsets of $J$ of Hausdorff dimension $1/2$ corresponding to arithmetic progressions of arbitrary lengths. As already illustrated in §$\S$1.2, an introduction of new ideas is necessary to prove our main results. Since the goal is to construct a subset of $E$, overlapping digits must be avoided. This means that elements of $A$ already appearing in the digit sequences of points in the seed set cannot be added. In order to avoid overlaps of digits, what we actually do in the second step is to add elements of $A\setminus S_{*}$. According to (1.8), the elements of $A$ must appear in the digit set of any point in $E_{S,A}$ with positive relative upper density equal to $\overline{d}(A|S)$. Hence, the upper density of $S_{*}$ relative to $S$ must be zero. This restriction requires an elaboration on the choice of $S_{*}$ in the first step. As such, the seed set we construct in §$\S$3.1 is of maximal Hausdorff dimension and least relative density of digits, called an extreme seed set.
Our main task in the second step is to add sufficiently many elements of $A\setminus S_{*}$ in appropriate positions to achieve the positive density $\overline{d}(A|S)$ without losing Hausdorff dimension of the seed set at all. This seemingly contradictory task can be accomplished by a sharp estimate of Hölder exponent of the associated elimination map, which essentially relies on the non-monotonicity of continued fraction digits for points in $E\setminus J$. See §$\S$3.2 for a general setup of the second step and relevant definitions, and §$\S$3.3 for the key estimate on the elimination map.
The rest of this paper consists of three sections and three appendices. In §$\S 2$ we summarize preliminary results needed for the proofs of the main results and Theorem 1.13. In §$\S 3$ we complete the first step of the construction outlined in §$\S$1.3, and then lay a ground work for the second step. In §$\S 4$ we perform the second step and complete the proofs of the main results. We also complete a proof of Theorem 1.13 in §$\S 4$. In Appendices A and B we prove Theorems 1.2 and 1.10 respectively. In Appendix C we leave comments on generalizations of our main results to a wide class of infinite Iterated Function Systems.
2. Preliminaries
This section summarizes preliminary results needed for the proofs of the main results and Theorem 1.13. In §$\S$2.1 we recall basic properties of continued fraction digits in the expansion (1.1). In §$\S$2.2 we introduce a Moran fractal, and provide two lower bounds of Hausdorff dimension of sets in $[0,1]$. In §$\S$2.3 we introduce a convergence exponent for infinite subsets of $\mathbb{N}$ and investigate its properties.
2.1. Basic properties of continued fraction digits
For each $n\in\mathbb{N},$ define non-negative integers $p_{n}$ and $q_{n}$ by the recursion formulas
$\begin{split}p_{-1}=1,\ p_{0}=0,\ p_{i}&=a_{i}p_{i-1}+p_{i-2}\ \text{ for }1\leq i\leq n,\\ q_{-1}=0,\ q_{0}=1,\ q_{i}&=a_{i}q_{i-1}+q_{i-2}\ \text{ for }1\leq i\leq n.\end{split}$ (2.1)
For $n\in\mathbb{N}$ and $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, we call the interval
$I(a_{1},\ldots,a_{n})=\begin{cases}\vspace{1mm}\left[\displaystyle{\frac{p_{n}}{q_{n}}},\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}\right)&\text{if}\ n\ \text{is even,}\\ \displaystyle{\left(\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}},\frac{p_{n}}{q_{n}}\right]}&\text{if}\ n\ \text{is odd,}\end{cases}$
an $n$-th fundamental interval. This interval represents the set of elements of $(0,1]$ which have the finite or infinite regular continued fraction expansion beginning by $a_{1},\ldots,a_{n}$, i.e.,
$I(a_{1},\ldots,a_{n})=\{x\in(0,1]\colon a_{i}(x)=a_{i}\ \text{ for }1\leq i\leq n\}.$
The next lemma estimates the length of each fundamental interval in terms of the set of digits prescribing it. Let ||$\cdot$|| ⋅ | denote the Lebesgue measure on $[0,1]$.
Lemma 2.1.
For every $n\in\mathbb{N}$ and every $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ we have
$\frac{1}{2}\prod_{i=1}^{n}\frac{1}{(a_{i}+1)^{2}}\leq|I(a_{1},\ldots,a_{n})|\leq\prod_{i=1}^{n}\frac{1}{a_{i}^{2}}.$
Proof.
By [16, 19], for every $n\in\mathbb{N}$ and every $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ we have
$|I(a_{1},\ldots,a_{n})|=\frac{1}{q_{n}(q_{n}+q_{n-1})}\ \text{ and }\ \prod_{i=1}^{n}a_{i}\leq q_{n}\leq\prod_{i=1}^{n}(a_{i}+1),$ (2.2)
where $q_{n-1}$, $q_{n}$ are given by (2.1). Since $q_{n-1}<q_{n}$ we have
$\frac{1}{2q_{n}^{2}}\leq|I(a_{1},\ldots,a_{n})|\leq\frac{1}{q_{n}^{2}}.$
Combining these with the double inequalities in (2.2) yields the desired ones. ∎
2.2. Lower bounds of Hausdorff dimension
Let $F=\bigcap_{n=0}^{\infty}E_{n}$ be a set in $[0,1]$ such that the following holds for every $n\geq 1$:
  • $E_{n-1}$ is the union of finitely many closed intervals of positive length, called $(n-1)$-th level intervals;
  • for each $(n-1)$-th level interval $I$, the number of $n$-th intervals contained in $E_{n}\cap I$ is at least $r_{n}$, and all of these $n$-th intervals are separated by gaps of length at least $\delta_{n}$, where $r_{n}\geq 2$ and $0<\delta_{n+1}<\delta_{n}$.
We call $F$ a Moran fractal with profile $(r_{n})_{n=1}^{\infty}$, $(\delta_{n})_{n=1}^{\infty}$. The Hausdorff dimension of a Moral fractal is estimated from below with its profile by the next lemma.
Lemma 2.2 ([8] pp.71–72).
Let $F\subset[0,1]$ be a Moran fractal with profile $(r_{n})_{n=1}^{\infty}$, $(\delta_{n})_{n=1}^{\infty}$. Then
$\dim_{\rm H}F\geq\liminf_{n\to\infty}\frac{\log(r_{1}r_{2}\cdots r_{n-1})}{-\log(r_{n}\delta_{n})}.$
Let $F\subset[0,1]$ be a set. We say $f\colon F\to[0,1]$ is Hölder continuous with exponent $\gamma\in(0,1]$ if there exists $C>0$ such that
$|f(x)-f(y)|\leq C|x-y|^{\gamma}\ \text{ for all }x,y\in F.$
Lemma 2.3 ([8] Proposition 3.3).
Let $F\subset[0,1]$ and let $f\colon F\to[0,1]$ be Hölder continuous with exponent $\gamma\in(0,1]$. Then $\dim_{\rm H}F\geq\gamma\cdot\dim_{\rm H}f(F).$
We say $f\colon F\to[0,1]$ is almost Lipschitz if for any $\gamma\in(0,1)$, $f$ is Hölder continuous with exponent $\gamma$. If $f$ is not Lipschitz continuous and almost Lipschitz, then the multiplicative constant $C$ grows to infinity as $\gamma$ increases to $1$, but this growth is inconsequential in the lower bound of Hausdorff dimension. Increasing $\gamma$ to $1$ in Lemma 2.3 we obtain the next lemma.
Lemma 2.4.
Let $F\subset[0,1]$ and let $f\colon F\to[0,1]$ be almost Lipschitz. Then
$\dim_{\rm H}F\geq\dim_{\rm H}f(F).$
2.3. Convergence exponent
For an infinite subset $S$ of $\mathbb{N}$, define
$\tau(S)=\inf\left\{t\geq 0\colon\sum_{a\in S}a^{-t}<\infty\right\}.$
We call $\tau(S)$ a convergence exponent of $S$. Note [^2] The convergence exponent was originally defined for infinite sequences of positive reals, see [30, §$\S$2]. Then the value ranges from 00 to $\infty$. that $\tau(S)\in[0,1]$. If we write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots$ then
$\tau(S)=\limsup_{n\to\infty}\frac{\log n}{\log a_{n}},$ (2.3)
see [30, §$\S$2]. The convergence exponent is often used in fractal dimension theory, see [15, 40] for example. The next lemma can be proved by a standard covering argument by fundamental intervals.
Lemma 2.5 ([15] Corollary 1).
For any infinite subset $S$ of $\mathbb{N}$, we have
$\dim_{\rm H}\{x\in(0,1)\setminus\mathbb{Q}\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\text{ and }a_{n}(x)\to\infty\text{ as }n\to\infty\}\leq\frac{\tau(S)}{2}.$
Remark 2.6.
The inequality in Lemma 2.5 is actually the equality for any infinite subset $S$ of $\mathbb{N}$ (see [40]), but we do not need this fact. The equality for $S=\mathbb{N}$ is well-known as Good’s theorem [13].
The next lemma links the convergence exponent to relative upper density. The case $A\neq S$ will only be used in the proof of Theorem 1.13.
Lemma 2.7.
If $S\subset\mathbb{N}$ has polynomial density with exponent $\alpha$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ we have
$\tau(A)=\frac{1}{\alpha}.$
Proof.
Write $A=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots$. Since $S$ has polynomial density with exponent $\alpha$, there exist $C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $n$ we have
$\frac{\log n}{\log a_{n}}=\frac{\log\#(A\cap[1,a_{n}])}{\log a_{n}}\leq\frac{\log(C_{2}/(\log a_{n})^{\beta})}{\log a_{n}}+\frac{1}{\alpha}.$
Letting $n\to\infty$ and combining the result with (2.3) we obtain $\tau(A)\leq 1/\alpha.$
In the case $A=S$, the same argument as above yields the reverse inequality $\tau(A)\geq 1/\alpha.$ In the case $A\neq S$, the same argument does not work since the positivity of $\overline{d}(A|S)$ does not imply that of $\limsup_{n\to\infty}\#(A\cap[1,a_{n}])/\#(S\cap[1,a_{n}])$. Hence we argue as follows. Take a strictly increasing sequence $(N_{k})_{k=1}^{\infty}$ of positive integers such that
$\lim_{k\to\infty}\frac{\#(A\cap[1,N_{k}])}{\#(S\cap[1,N_{k}])}=\overline{d}(A|S)>0.$
Let $\varepsilon\in(0,1/\alpha)$. Since $S$ has polynomial density with exponent $\alpha$, there exist $C_{1}>0$ and $\beta\geq 0$ such for all sufficiently large $k$ we have
#$\begin{split}\frac{\#(A\cap[N_{k}^{\varepsilon},N_{k}])}{\#(S\cap[1,N_{k}])}&\geq\frac{\#(A\cap[1,N_{k}])}{\#(S\cap[1,N_{k}])}-\frac{\#(A\cap[1,N_{k}^{\varepsilon}])}{\#(S\cap[1,N_{k}])}\\ &\geq\frac{\overline{d}(A|S)}{2}-\frac{(\log N_{k})^{\beta}}{C_{1}N_{k}^{1/\alpha-\varepsilon}}\geq\frac{\overline{d}(A|S)}{3}.\end{split}$
Rearranging this gives
#$\#(A\cap[N_{k}^{\varepsilon},N_{k}])\geq\frac{\overline{d}(A|S)}{3}\#(S\cap[1,N_{k}])\geq\frac{\overline{d}(A|S)}{3}\frac{C_{1}N_{k}^{1/\alpha}}{(\log N_{k})^{\beta}}.$
Let $0<t<1/\alpha$. For all sufficiently large $k\geq 1$ we have
$\sum_{a\in A\cap[N_{k}^{\varepsilon},N_{k}]}a^{-t}>N_{k}^{-t}\#(A\cap[N_{k}^{\varepsilon},N_{k}])\geq\frac{\overline{d}(A|S)}{3}\frac{C_{1}N_{k}^{1/\alpha-t}}{(\log N_{k})^{\beta}},$
Letting $k\to\infty$ we conclude that $\sum_{a\in A}a^{-t}=\infty$, namely $\tau(A)\geq t$. Increasing $t$ to $1/\alpha$ we obtain $\tau(A)\geq 1/\alpha$.∎
The next result obtained in the multifractal analysis of convergence exponent for continued fractions [9] is a key ingredient in the proof of Theorem 1.13.
Lemma 2.8 ([9], Theorem 2.2).
We have
[…]
3. Key constructions
This section is dedicated to key constructions in the proofs of the main results. In §$\S$3.1 we introduce and construct an extreme seed set. In §$\S$3.2 we describe a procedure of constructing a new Moran-like fractal of $(0,1)\setminus\mathbb{Q}$ by adding digits into a given seed set. In §$\S$3.3 we prove that an elimination map associated with the construction in §$\S$3.2 is almost Lipschitz.
3.1. Construction of extreme seed sets
Let $S\subset\mathbb{N}$ have polynomial density. For an infinite subset $S_{*}$ of $S$ and positive reals $t,L$ satisfying $t>L>1$, define
$R_{t,L}(S_{*})=\{x\in E\colon a_{n}(x)\in S_{*}\cap[t^{n},Lt^{n}]\ \text{ for every}\ n\geq 1\}.$
Non-empty sets of this form are called seed sets. Seed sets are Moran-like fractals, and their Hausdorff dimensions do not exceed $1/(2\alpha)$ by Lemmas 2.5 and 2.7. In fact, one can decrease the relative upper density of $S_{*}$ without decreasing the Hausdorff dimension of the corresponding seed set. We say the seed set $R_{t,L}(S_{*})$ is an extreme seed set if $S_{*}$ does not contain consecutive integers, and
$\overline{d}(S_{*}|S)=0\ \text{ and }\ \dim_{\rm H}R_{t,L}(S_{*})=\frac{1}{2\alpha}.$
Extreme seed sets are of maximal Hausdorff dimension and least relative density of digits. The assumption of not containing consecutive integers is not essential, but a little facilitates the proof of Proposition 3.5 in §$\S$3.2.
Proposition 3.1.
If $S\subset\mathbb{N}$ has polynomial density, then there exist $S_{*}\subset S$ and positive reals $t,L$ such that $t>L>1$, $t>2$ and $R_{t,L}(S_{*})$ is an extreme seed set.
For a proof of this proposition we need two preliminary lemmas. First, define a function $\nu\colon[1,\infty)\to\mathbb{N}$ by
$\nu(\xi)=k\ \text{ for }\xi\in[k!,(k+1)!),\ k\in\mathbb{N}.$ (3.1)
This is a slow increasing function that has the following upper bound.
Lemma 3.2.
For any $t>e$ there exists $\lambda>1$ such that
$\nu(t^{n})\leq(n\log t)^{\lambda}\ \text{ for every }n\in\mathbb{N}.$
Proof.
We claim that $\nu(t^{n})\leq n\log t$ for every $n\in\mathbb{N}$ but finitely many ones. For otherwise the definition of $\nu$ in (3.1) would yield $t^{n}>\lfloor n\log t\rfloor!$ for infinitely many $n$, which is a contradiction. There exists a constant $c(t)\geq 1$ such that $\nu(t^{n})\leq c(t)n\log t$ for every $n\in\mathbb{N}$. Take $\lambda>1$ such that $c(t)\leq(\log t)^{\lambda-1}$. ∎
The next general lemma allows us to extract from an arbitrary infinite subset of $\mathbb{N}$ a set of zero relative density with definite growth bounds.
Lemma 3.3.
Let $S$ be an infinite subset of $\mathbb{N}$ and write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots$. There exists $S_{*}\subset S$ that does not contain consecutive integers and satisfies
$\frac{1}{2\nu(\Delta)}\leq\frac{\#\{1\leq n\leq\Delta\colon a_{n}\in S_{*}\}}{\Delta}\leq\frac{3}{\nu(\Delta)}$ (3.2)
for all sufficiently large $\Delta\geq 1$. In particular, $\overline{d}(S_{*}|S)=0$.
Proof.
For each $k\in\mathbb{N}$, taking every $k$ elements of $\mathbb{N}$ with respect to the natural order one can construct a subset of $\mathbb{N}$ with density $1/(k+1)$. Our construction here relies on this simple observation. Let
$Q=\bigcup_{k=2}^{\infty}\{k!+ik\colon i=0,\ldots,k!-1\}\subset\mathbb{N}.$
The following interpretation of $Q$ should be useful along the way. Split $[2,\infty)=\bigcup_{k=2}^{\infty}[k!,(k+1)!)$. For each integer $k\geq 2$, further split $[k!,(k+1)!)$ into consecutive left-closed, right-open subintervals containing exactly $k$ integers:
$[k!,(k+1)!)=[k!,k!+k)\cup[k!+k,k!+2k)\cup\cdots\cup[k!+(k!-1)k,(k+1)!).$
Then $Q$ consists of the left endpoints of all these subintervals. Note that $Q$ does not contain consecutive integers. Set
$S_{*}=\{a_{n}\in S\colon n\in Q\}.$
Then $S_{*}$ does not contain consecutive integers.
For every $\Delta\geq 1$ we have
#$\#\{1\leq n\leq\Delta\colon a_{n}\in S_{*}\}=\#(Q\cap[1,\Delta]).$
If $\Delta\geq 1$ is sufficiently then we have
#$\begin{split}\#(Q\cap[1,\Delta])&=\#(Q\cap[1,(k-1)!))+\#(Q\cap[(k-1)!,k!))+\#(Q\cap[k!,\Delta])\\ &\leq\#(Q\cap[1,(k-1)!))+\frac{\#([(k-1)!,k!)\cap\mathbb{N})}{k-1}+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}\\ &\leq(k-1)!+(k-1)!+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}\leq\frac{3\Delta}{\nu(\Delta)},\end{split}$
and
#$\begin{split}\#(Q\cap[1,\Delta])&\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{i}+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}-1\\ &\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{k}+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}-1\\ &=\frac{\#([2,\Delta]\cap\mathbb{N})}{k}-1\geq\frac{\Delta}{2\nu(\Delta)},\end{split}$
where $k\geq 1$ is the integer with $\Delta\in[k!,(k+1)!)$. This verifies (3.2).
The ordering of elements of $S$ implies #$\#(S_{*}\cap[1,N])=\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}$ for every $N\geq 1$ with $S\cap[1,N]\neq\emptyset$. By the second inequality in (3.2), for all sufficiently large $N\geq 1$ we have
#$\frac{\#(S_{*}\cap[1,N])}{\#(S\cap[1,N])}=\frac{\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}}{\#(S\cap[1,N])}\leq\frac{3}{\nu(\#(S\cap[1,N]))}.$
Letting $N\to\infty$ we obtain $\overline{d}(S_{*}|S)=0$ as required. ∎
Proof of Proposition 3.1.
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha$ and write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots.$ Let $S_{*}$ be a subset of $S$ for which the conclusion of Lemma 3.3 holds. It suffices to show that $\dim_{\rm H}R_{t,L}(S_{*})\geq 1/(2\alpha)$ for some $t,L$ with $t>L>1$ and $t>2$.
By the ordering of elements of $S$, for any $N\geq 1$ with $S\cap[1,N]\neq\emptyset$ we have #$\#(S_{*}\cap[1,N])=\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}$. Moreover (3.2) gives
#$\begin{split}\frac{\#(S\cap[1,N])}{2\nu(\#(S\cap[1,N]))}\leq\#(S_{*}\cap[1,N])&\leq\frac{3\#(S\cap[1,N])}{\nu(\#(S\cap[1,N]))}.\end{split}$ (3.3)
Let $C_{1}>0$, $C_{2}>0$, $\beta\geq 0$ be the constants in the definition of polynomial density of $S$ with exponent $\alpha$. Set
$L=(12C_{1}^{-1}C_{2}2^{\beta})^{\alpha}>1.$ (3.4)
Let $t>\max\{L,e\}$ satisfy
$\nu(\#(S\cap[1,t]))\geq 2.$ (3.5)
Since $t>L$, if $t$ is sufficiently large then for every $n\in\mathbb{N}$ we have
$C_{1}C_{2}^{-1}2^{-\beta}L^{-1/\alpha}\leq\frac{\#(S\cap[1,t^{n}])}{\#(S\cap[1,Lt^{n}])}\leq C_{1}^{-1}C_{2}2^{\beta}L^{-1/\alpha}.$ (3.6)
For each $n\in\mathbb{N}$, fix $k=k(n)\in\mathbb{N}$ such that $k!\leq\#(S\cap[1,t^{n}])<(k+1)!.$ If $n$ is sufficiently large so that $k+2\geq(C_{1}C_{2}^{-1}2^{-\beta}L^{-1/\alpha})^{-1}$, then
$k!\leq\#(S\cap[1,t^{n}])\leq\#(S\cap[1,Lt^{n}])\leq\frac{\#(S\cap[1,t^{n}])}{C_{1}C_{2}^{-1}2^{-\beta}L^{-1/\alpha}}<(k+2)!.$
By the definition of $\nu$ in (3.1), this estimate implies
$0\leq\nu(\#(S\cap[1,Lt^{n}]))-\nu(\#(S\cap[1,t^{n}]))\leq 1$ (3.7)
By $t>e$ and Lemma 3.2, there exists $\lambda>1$ such that for every $n\in\mathbb{N}$,
$\nu(\#(S\cap[1,t^{n}]))\leq\nu(t^{n})\leq C_{1}(n\log t)^{\lambda}.$ (3.8)
Using (3.3), (3.5) and then (3.6), (3.7), for each integer $n\geq 1$ we have
#$\begin{split}\#(S_{*}\cap[t^{n},Lt^{n}])&\geq\#(S_{*}\cap[1,Lt^{n}])-\#(S_{*}\cap[1,t^{n}])\\ &\geq\frac{\#(S\cap[1,Lt^{n}])}{2\nu(\#(S\cap[1,Lt^{n}]))}-\frac{3\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}\\ &\geq\frac{C_{1}C_{2}^{-1}2^{-\beta}L^{1/\alpha}\#(S\cap[1,t^{n}])}{2(\nu(\#(S\cap[1,t^{n}])+1)}-\frac{3\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}.\end{split}$
By (3.5), the last expression is bounded from below by
#$\#(S\cap[1,t^{n}])\left(\frac{C_{1}C_{2}^{-1}2^{-\beta}L^{1/\alpha}}{3\nu(\#(S\cap[1,t^{n}]))}-\frac{3}{\nu(\#(S\cap[1,t^{n}]))}\right),$
which equals #$\#(S\cap[1,t^{n}])/\nu(\#(S\cap[1,t^{n}]))$ by (3.4). Set $\rho=\beta+\lambda$. Using (3.8) and the assumption that $S$ has polynomial density, we obtain
#$\begin{split}\#(S_{*}\cap[t^{n},Lt^{n}])\geq\frac{\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}&\geq C_{1}\frac{t^{\frac{n}{\alpha}}}{(n\log t)^{\beta}}\cdot\frac{1}{C_{1}(n\log t)^{\lambda}}\\ &=\frac{t^{\frac{n}{\alpha}}}{(n\log t)^{\rho}}.\end{split}$ (3.9)
We now set
$F=\bigcap_{n=1}^{\infty}\bigcup_{\begin{subarray}{c}a_{i}\in{S_{*}\cap[t^{i},Lt^{i}]}\\ 1\leq i\leq n\end{subarray}}\overline{I(a_{1},\ldots,a_{n})},$
where the bar denotes the closure. Since $t>L$, the intervals $[t^{n},Lt^{n}]$, $n\in\mathbb{N}$ are pairwise disjoint. The Moran fractal $F$ is contained in $R_{t,L}(S_{*})$ up to countable sets consisting of the boundary points of the level intervals. For every $n\geq 2$ and every $(n-1)$-th level interval of $F$, the number of $n$-th level intervals contained in that interval is at least ${\lfloor t^{\frac{n}{\alpha}}/(n\log t)^{\rho}\rfloor}$ by (3.9). Since $S_{*}$ does not contain consecutive integers, these $n$-th level intervals are separated at least by some $n$-th fundamental intervals. Hence, the gap between the $n$-th level intervals is at least
$\min\{|I(a_{1},\ldots,a_{n})|\colon t^{i}\leq a_{i}\leq Lt^{i},1\leq i\leq n\}.$
By Lemma 2.1 and $t>L$, we have
$|I(a_{1},\ldots,a_{n})|\geq\frac{1}{2}\prod_{i=1}^{n}\frac{1}{(a_{i}+1)^{2}}\geq\frac{1}{2}\prod_{i=1}^{n}\frac{1}{(Lt^{i}+1)^{2}}\geq\frac{1}{2}\prod_{i=1}^{n}\frac{1}{(Lt^{i+1})^{2}}.$
So, $F$ is a Moran fractal with profile $({\lfloor t^{\frac{n}{\alpha}}/(n\log t)^{\rho}\rfloor})_{n=1}^{\infty}$, $((1/2)\prod_{i=1}^{n}(Lt^{i+1})^{-2})_{n=1}^{\infty}$. Lemma 2.2 gives
$\dim_{\rm H}F\geq\liminf_{n\to\infty}\frac{\log(2\prod_{i=1}^{n-1}{\lfloor t^{\frac{i}{\alpha}}/(i\log t)^{\rho}\rfloor})}{-\log\left((\lfloor t^{\frac{n}{\alpha}}\rfloor/(n\log t)^{\rho})\prod_{i=1}^{n}(Lt^{i+1})^{-2}\right)}.$
Let us agree that for two eventually positive functions $f$, $g$ on $\mathbb{N}$, the expression $f(n)\sim g(n)$ indicates that $f(n)/g(n)\to 1$ as $n\to\infty$. Direct calculations show
$\log\left(2\prod_{i=1}^{n-1}{\left\lfloor\frac{t^{\frac{i}{\alpha}}}{(i\log t)^{\rho}}\right\rfloor}\right)$ $=\log 2+\sum_{i=1}^{n-1}\log\left\lfloor\frac{t^{\frac{i}{\alpha}}}{(i\log t)^{\rho}}\right\rfloor$
$\sim\log 2+\sum_{i=1}^{n-1}\left(\log t^{\frac{i}{\alpha}}-\log(i\log t)^{\rho}\right)\sim\frac{1}{2\alpha}n^{2}\log t,$
and
$\begin{split}-\log\left(\frac{\lfloor t^{\frac{n}{\alpha}}\rfloor}{(n\log t)^{\rho}}\prod_{i=1}^{n}\frac{1}{(Lt^{i+1})^{2}}\right)&{\sim}-\frac{n}{\alpha}\log t+\rho\log(n\log t)+\sum_{i=1}^{n}2\log(Lt^{i+1})\\ &\sim n^{2}\log t.\end{split}$
Combining these two expressions yields $\dim_{\rm H}R_{t,L}(S_{*})\geq\dim_{\rm H}F\geq 1/(2\alpha)$. This completes the proof of Proposition 3.1. ∎
Corollary 3.4.
If $S\subset\mathbb{N}$ has polynomial density with exponent $\alpha$, then
$\dim_{\rm H}\{x\in J\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}=\dim_{\rm H}\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}=\frac{1}{2\alpha}.$
Proof.
On the one hand, by Lemmas 2.5 and 2.7 we have
$\dim_{\rm H}\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}\leq\frac{1}{2\alpha}.$
On the other hand, Proposition 3.1 implies
$\dim_{\rm H}\{x\in J\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}\geq\frac{1}{2\alpha}.$
Since $J\subset E$, these two inequalities yield the desired equalities. ∎
3.2. Adding digits into seed sets
Let $A$, $B$ be infinite subsets of $\mathbb{N}$. Let $t>L>1$ and suppose $R_{t,L}(B)\neq\emptyset$. We add elements of $A$ into the digit sequences of points in the seed set $R_{t,L}(B)$, and construct a new subset of $(0,1)\setminus\mathbb{Q}$ in the following manner.
Let $M_{0}=0$, and let $(M_{k})_{k=1}^{\infty}$ be a strictly increasing sequence of positive integers. Let $(W_{k})_{k=1}^{\infty}$ be a sequence of finite subsets of $A$. Let $y\in R_{t,L}(B)$ and put $a_{0}(y)=0$ for convenience. For each $k\geq 1$ with $W_{k}\neq\emptyset$, we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ in the increasing order [^3] This ordering in $W_{k}$ is not essential. For a related comment, see the footnote in Corollary 1.3. to define a new sequence
$\begin{split}&\ldots,a_{M_{k}-1}(y),\ a_{M_{k}}(y),\ w_{1}^{(k)},\ldots,w_{i(k)}^{(k)},\ a_{M_{k}+1}(y),\ldots\\ &\ldots,a_{M_{k+1}}(y),\ w_{1}^{(k+1)},\ldots,w_{i(k+1)}^{(k+1)},\ a_{M_{k+1}+1}(y),\ldots\end{split}$
where $W_{k}=\{w^{(k)}_{i}\colon i=1,\ldots,i(k)\}$, $w^{(k)}_{1}<\cdots<w^{(k)}_{i(k)}$. Let $x(y)$ denote the point in $(0,1)\setminus\mathbb{Q}$ whose regular continued fraction expansion (1.1) is given by this new sequence. Let $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ denote the collection of these points:
$R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})=\{x(y)\in(0,1)\setminus\mathbb{Q}\colon y\in R_{t,L}(B)\}.$
The map $y\in R_{t,L}(B)\mapsto x(y)\in R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ is bijective. Let
$f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}\colon R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})\to R_{t,L}(B)$
denote the inverse of this map. We call $f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}$ an elimination map. It eliminates all the digits in $\bigcup_{k=1}^{\infty}W_{k}$. Clearly, if $A\cap B=\emptyset$ and the elements of $(W_{k})_{k=1}^{\infty}$ are pairwise disjoint then $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ is contained in $E$.
3.3. Almost Lipschitzness of the elimination map
We need to deduce a lower bound of the Hausdorff dimension of the set $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ constructed in §$\S$3.2. This can be done by evaluating the Hölder exponent of the elimination map. This map is merely Hölder in general and as a result the Hausdorff dimension of $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ may drop from that of $R_{t,L}(B)$. However, under appropriate choices of $(M_{k})_{k=1}^{\infty}$ and $(W_{k})_{k=1}^{\infty}$, the Hölder exponent can be made closer and closer to $1$ on deeper levels of $R_{t,L}(B)$, so that the elimination map becomes almost Lipschitz, as stated in the proposition below.
Proposition 3.5.
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a monotone decreasing sequence of positive reals converging to 00. Let $A$, $B$ be infinite subsets of $\mathbb{N}$ and suppose $B$ does not contain consecutive integers. Let $t>L>1$, $t>2$ and suppose $R_{t,L}(B)\neq\emptyset$. Let $M_{0}=0$ and let $(M_{k})_{k=1}^{\infty}$ be a strictly increasing sequence of positive integers, and let $(W_{k})_{k=1}^{\infty}$ be a sequence of finite subsets of $A$ such that for every $k\geq 1$,
$\prod_{i\in W_{k}}\frac{1}{(i+1)^{2}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ if }W_{k}\neq\emptyset$ (3.10)
and
$\frac{1}{2}\prod_{i=1}^{n+1}\frac{1}{(t^{i+2})^{2}}\cdot\left(\prod_{i=1}^{n}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\geq\left(\prod_{i=1}^{n}\frac{1}{t^{2i}}\right)^{1+4\varepsilon_{k}}\ \text{ for every }n\geq M_{k}.$ (3.11)
Then $f=f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}$ is almost Lipschitz.
Proof.
A key observation motivating the proof of this proposition is that the influence of adding the elements of $W_{k}$ is absorbed into the contribution from the digits $a_{M_{k-1}+1}(y),\ldots,a_{M_{k}}(y)$, $y\in R_{t,L}(B)$ by virtue of (3.10). The role of (3.11) is auxiliary, to be used after (3.10).
Let us abbreviate $R_{t,L}(B)$, $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ to $R(B)$, $R(B,A)$ respectively. Let $k\in\mathbb{N}$. For a pair $y_{1},y_{2}$ of distinct points in $R(B)$, let $s(y_{1},y_{2})$ denote the minimal integer $n\geq 0$ such that $a_{n+1}(y_{1})\neq a_{n+1}(y_{2}).$ Take an integer $N\geq M_{k}$ such that for any pair $y_{1},y_{2}$ of distinct points in $R(B)$ with $s(y_{1},y_{2})<M_{k}$, there exists $i\in\{1,\ldots,N\}$ such that $a_{i}(f^{-1}(y_{1}))\neq a_{i}(f^{-1}(y_{2}))$. Since $\sup_{x\in R(B,A)}\max\{a_{1}(x),\ldots,a_{N}(x)\}$ is finite, there is a constant $C>0$ depending on $N$ such that for points $x_{1},x_{2}$ in $R(B,A)$ satisfying $a_{i}(x_{1})\neq a_{i}(x_{2})$ for some $i\leq N$, we have
$|f(x_{1})-f(x_{2})|\leq 1\leq C|x_{1}-x_{2}|.$ (3.12)
Let $x_{1},x_{2}\in R(B,A)$ be distinct and put $y_{1}=f(x_{1})$, $y_{2}=f(x_{2})$. If $s(y_{1},y_{2})<M_{k}$ then (3.12) gives $|y_{1}-y_{2}|\leq C|x_{1}-x_{2}|$. Suppose $s(y_{1},y_{2})=n\geq M_{k}$. Without loss of generality we may assume $a_{n+1}(y_{1})<a_{n+1}(y_{2}).$ For each $q\in\mathbb{N}$, set $m_{q}=\#W_{q}$. Let $q\geq k$ be the integer such that $M_{q}\leq n<M_{q+1}.$ The definition of the elimination map $f$ implies
$a_{i}(x_{1})=a_{i}(x_{2})\ \text{ for }1\leq i\leq n+m_{1}+\cdots+m_{q},$
and
$a_{n+m_{1}+\cdots+m_{q}+1}(x_{1})=a_{n+1}(y_{1})<a_{n+1}(y_{2})=a_{n+m_{1}+\cdots+m_{q}+1}(x_{2}).$
Since $x_{1}$ and $x_{2}$ are contained in $I(a_{1}(x_{1}),\ldots,a_{n+m_{1}+\cdots+m_{q}}(x_{1}))$ and $B$ does not contain consecutive integers, we have $a_{n+1}(y_{2})-a_{n+1}(y_{1})\geq 2$. Then $x_{1}$ and $x_{2}$ are separated by the fundamental interval
$I(a_{1}(x_{1}),\ldots,a_{n+m_{1}+\cdots+m_{q}}(x_{1}),a_{n+m_{1}+\cdots+m_{q}+1}(x_{1})+1).$
We estimate $|x_{1}-x_{2}|$ from below and $|y_{1}-y_{2}|$ from above. Using the first inequality in Lemma 2.1 and $a_{i}(y_{1})\leq Lt^{i}$ for $11\leq i\leq n+$ 1 we have
$\begin{split}|x_{1}-x_{2}|&\geq|I(a_{1}(x_{1}),\ldots,a_{n+m_{1}+\cdots+m_{q}}(x_{1}),a_{n+m_{1}+\cdots+m_{q}+1}(x_{1})+1)|\\ &\geq\frac{1}{2\left((a_{1}(x_{1})+1)\cdots(a_{n+m_{1}+\cdots+m_{q}}(x_{1})+1)(a_{n+m_{1}+\cdots+m_{q}+1}(x_{1})+1)\right)^{2}}\\ &=\frac{1}{2}\prod_{i=1}^{n+1}\frac{1}{(a_{i}(y_{1})+1)^{2}}\cdot\prod_{j=1}^{q}\prod_{i\in W_{j}}\frac{1}{(i+1)^{2}}\\ &\geq\frac{1}{2}\prod_{i=1}^{n+1}\frac{1}{(Lt^{i}+1)^{2}}\cdot\prod_{j=1}^{q}\left(\prod_{i=M_{j-1}+1}^{M_{j}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{j}}\\ &=\frac{D_{k}}{2}\prod_{i=1}^{n+1}\frac{1}{(Lt^{i}+1)^{2}}\cdot\prod_{j=k+1}^{q}\left(\prod_{i=M_{j-1}+1}^{M_{j}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{j}},\end{split}$
where
$D_{k}=\prod_{j=1}^{k}\left(\prod_{i=M_{j-1}+1}^{M_{j}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{j}}.$
To deduce the last inequality we have used (3.10). Using $t>L$ and $t>2$ further to bound the last expression, we obtain
$\begin{split}|x_{1}-x_{2}|&\geq\frac{D_{k}}{2}\prod_{i=1}^{n+1}\frac{1}{(Lt^{i}+1)^{2}}\cdot\prod_{j=1}^{q}\left(\prod_{i=M_{j-1}+1}^{M_{j}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\\ &\geq\frac{D_{k}}{2}\prod_{i=1}^{n+1}\frac{1}{(t^{i+2})^{2}}\cdot\left(\prod_{i=1}^{n}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}.\end{split}$ (3.13)
Since $s(y_{1},y_{2})=n\geq M_{k}$, (3.11) and (3.13) yield
$|x_{1}-x_{2}|\geq D_{k}\left(\prod_{i=1}^{n}\frac{1}{t^{2i}}\right)^{1+4\varepsilon_{k}}.$ (3.14)
The second inequality in Lemma 2.1 gives
$\prod_{i=1}^{n}\frac{1}{t^{2i}}\geq\prod_{i=1}^{n}\frac{1}{a_{i}(y_{1})^{2}}\geq|I(a_{1}(y_{1}),\ldots,a_{n}(y_{1}))|\geq|y_{1}-y_{2}|.$ (3.15)
Combining (3.14) and (3.15) we obtain
$|x_{1}-x_{2}|\geq D_{k}|y_{1}-y_{2}|^{1+4\varepsilon_{k}}.$
Setting $C_{k}=D_{k}^{\frac{1}{1+4\varepsilon_{k}}}$ yields $|f(x_{1})-f(x_{2})|\leq C_{k}|x_{1}-x_{2}|^{\frac{1}{1+4\varepsilon_{k}}}$. Since $\varepsilon_{k}\to 0$ as $k\to\infty$, $f$ is almost Lipschitz. ∎
4. Proofs of the main results
In this last section we complete the proofs of the main results. Since Theorem 1.1(a) follows from Theorem 1.4, we prove Theorem 1.4 in §$\S$4.1, and then prove Theorem 1.1(b) in §$\S$4.2. Finally we prove Theorem 1.13 in §$\S$4.3.
4.1. Proof of Theorem 1.4
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha$. By Proposition 3.1 there exist $S_{*}\subset S$ not containing consecutive integers and positive reals $t$, $L$ with $t>L>1$ and $t>2$ such that $R_{t,L}(S_{*})$ is an extreme seed set of Hausdorff dimension $1/(2\alpha)$. Let $A\subset S$ satisfy $\overline{d}(A|S)>0$. We have
$\overline{d}((A\setminus S_{*})|S)=\overline{d}(A|S).$ (4.1)
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a monotone decreasing sequence of positive reals converging to 00. We set $M_{0}=0$ for convenience, and choose a strictly increasing sequence $(M_{k})_{k=1}^{\infty}$ of positive integers greater than one such that all the following conditions hold:
$\lim_{k\to\infty}\frac{\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap[1,\sqrt{M_{k}}])}=\overline{d}(A\setminus S_{*}|S);$ (4.2)
$\lim_{k\to\infty}\frac{t^{M_{k-1}}}{\sqrt{M_{k}}^{1/\alpha}}=0;$ (4.3)
$\ Lt^{M_{k-1}}+\sqrt{M_{k}}<t^{M_{k}}\ \text{ for every }k\geq 1;$ (4.4)
$\frac{1}{(\sqrt{M_{k}}+Lt^{M_{k-1}}+1)^{2\sqrt{M_{k}}}}\geq\frac{1}{t^{2\varepsilon_{k}M_{k}^{2}}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1;$ (4.5)
$\inf_{n\geq M_{k}}\prod_{i=1}^{n+1}\frac{1}{(t^{i+2})^{2}}\cdot\left(\prod_{i=1}^{n}\frac{1}{t^{2i}}\right)^{-1-\varepsilon_{k}}\geq 2\ \text{ for every }k\geq 1.$ (4.6)
Clearly (4.3), (4.4), (4.5) hold if $M_{k}$ is chosen to be sufficiently large compared to $M_{k-1}$. The role of (4.6) is auxiliary, which holds if $M_{k}$ is sufficiently large depending on $\varepsilon_{k}$.
For each $k\in\mathbb{N}$ we put
$I_{k}=[Lt^{M_{k-1}}+1,Lt^{M_{k-1}}+\sqrt{M_{k}}]$
and
$W_{k}=(A\setminus S_{*})\cap I_{k}.$
Taking a subsequence of $(M_{k})_{k=1}^{\infty}$ if necessary we may assume
$\lim_{k\to\infty}\frac{\min I_{k}}{\max I_{k}}=0.$ (4.7)
Since $S$ has polynomial density, it intersects both $I_{k}$ and $[1,\sqrt{M_{k}}]$ for all sufficiently large $k$. In the next two paragraphs we show that the two differences
#$\frac{\#(A\cap I_{k})}{\#(S\cap I_{k})}-\frac{\#W_{k}}{\#(S\cap I_{k})}=\frac{\#(S_{*}\cap I_{k})}{\#(S\cap I_{k})}$
and
|#$\begin{split}\left|\frac{\#W_{k}}{\#(S\cap I_{k})}-\frac{\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap[1,\sqrt{M_{k}}])}\right|\leq&\left|\frac{\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap I_{k})}\right|\\ &+\left|\frac{\#(S\cap[1,\sqrt{M_{k}}])-\#(S\cap I_{k})}{\#(S\cap I_{k})}\right|\end{split}$
converge to 00 as $k\to\infty$.
Since $S$ has polynomial decay with exponent $\alpha\geq 1$, there exist constants $C_{1}>0$ and $\beta\geq 0$ such that
$\begin{split}&\frac{1}{\#(S\cap I_{k})}\leq\left(\frac{C_{1}(Lt^{M_{k-1}}+\sqrt{M_{k}})^{1/\alpha}}{(\log(Lt^{M_{k-1}}+\sqrt{M_{k}}))^{\beta}}-Lt^{M_{k-1}}\right)^{-1}.\end{split}$
Since $S$ has polynomial decay with exponent $\alpha\geq 1$ and $\overline{d}(S_{*}|S)=0$, there exist $C_{2}>0$ and for any $\varepsilon>0$ there exists $k_{0}\geq 1$ such that for every $k\geq k_{0}$,
#$\#(S_{*}\cap I_{k})\leq\#(S_{*}\cap[1,Lt^{M_{k-1}}+\sqrt{M_{k}}]\leq\varepsilon\cdot\frac{C_{2}(Lt^{M_{k-1}}+\sqrt{M_{k}})^{1/\alpha}}{(\log(Lt^{M_{k-1}}+\sqrt{M_{k}}))^{\beta}}.$
Combining these two inequalities and then using (4.3), we obtain
$\lim_{k\to\infty}\left|\frac{\#(A\cap I_{k})}{\#(S\cap I_{k})}-\frac{\#W_{k}}{\#(S\cap I_{k})}\right|=0,$ (4.8)
as required.
Moving on to the second difference, for the numerator of the first fraction in the right-hand side we have
|#$\begin{split}&|\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])|\\ =&|\#((A\setminus S_{*})\cap[Lt^{M_{k-1}}+1,\sqrt{M_{k}}])+\#((A\setminus S_{*})\cap[\sqrt{M_{k}}+1,\sqrt{M_{k}}+Lt^{M_{k-1}}])\\ &-\#((A\setminus S_{*})\cap[1,Lt^{M_{k-1}}])-\#((A\setminus S_{*})\cap[Lt^{M_{k-1}}+1,\sqrt{M_{k}}])|\leq Lt^{M_{k-1}}.\end{split}$
For the numerator of the second fraction, a similar calculation shows
|#$|\#(S\cap[1,\sqrt{M_{k}}])-\#(S\cap I_{k})|\leq Lt^{M_{k-1}}.$
Condition (4.3) implies
$\lim_{k\to\infty}\frac{\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap I_{k})}=0$
and
$\lim_{k\to\infty}\frac{\#(S\cap[1,\sqrt{M_{k}}])-\#(S\cap I_{k})}{\#(S\cap I_{k})}=0.$
Hence we obtain
$\lim_{k\to\infty}\left|\frac{\#W_{k}}{\#(S\cap I_{k})}-\frac{\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap[1,\sqrt{M_{k}}])}\right|=0,$ (4.9)
as required.
Combining (4.8), (4.9) and then using (4.1), (4.2),
$\begin{split}\lim_{k\to\infty}\frac{\#(A\cap I_{k})}{\#(S\cap I_{k})}&=\lim_{k\to\infty}\frac{\#W_{k}}{\#(S\cap I_{k})}=\overline{d}(A|S)>0.\end{split}$ (4.10)
This implies that $W_{k}$ is non-empty for all sufficiently large $k$. Since $\max I_{k}<\min I_{k+1}$ for every $k\geq 1$ by (4.4), $\{W_{k}\colon k\in\mathbb{N}\}$ consists of pairwise disjoint, finite subsets of $A\setminus S_{*}$. We set
$E_{S,A}=R_{t,L}(S_{*},A\setminus S_{*},(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}).$
The construction of this set was described in §$\S$3.2. For each $y\in R_{t,L}(S_{*})$ we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ at position $M_{k}$ to define a new sequence. Then $E_{S,A}$ is the set of points whose regular continued fraction expansions are given by these new sequences.
We claim that $E_{S,A}\subset E$. Indeed, let $x\in E_{S,A}$ and let $y$ be the corresponding point in $R_{t,L}(S_{*})$, namely $y=f(x)$. Since $(a_{n}(y))_{n=1}^{\infty}$ is a strictly increasing sequence in $S_{*}$ and $\bigsqcup_{k=1}^{\infty}W_{k}\subset A\setminus S_{*}$, for all $m$, $n\in\mathbb{N}$ distinct we have $a_{m}(x)\neq a_{n}(x)$, namely $x\in E$. By the construction of $E_{S,A}$, we have
#$\#W_{k}\leq\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}\leq\#(A\cap I_{k}),$
and (4.10) implies
$\lim_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}}{\#(S\cap I_{k})}=\overline{d}(A|S).$ (4.11)
Since (3.10) and (3.11) follow from (4.5) and (4.6) respectively, the corresponding elimination map is almost Lipschitz by Proposition 3.5. Lemma 2.4 gives
$\dim_{\rm H}E_{S,A}\geq\dim_{\rm H}R_{t,L}(S_{*})=\frac{1}{2\alpha}.$ (4.12)
It is clear from the construction that $E_{S,A}\subset\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$. Combining (4.12) and Corollary 3.4 we obtain (1.6).
Since $S$ has polynomial density with exponent $\alpha$, there exist $C_{1},C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $k\geq 1$ we have
#$\frac{\#(S\cap[1,\min I_{k}))}{\#(S\cap[1,\max I_{k}])}\leq C_{1}C_{2}^{-1}\frac{(\min I_{k})^{1/\alpha}}{(\max I_{k})^{1/\alpha}}\frac{(\log(\max I_{k}))^{\beta}}{(\log(\min I_{k}))^{\beta}}.$
On the right-hand side, using $\lim_{k\to\infty}\min I_{k}=\infty$ and (4.7) we have
$\lim_{k\to\infty}\frac{(\min I_{k})^{1/\alpha}}{(\max I_{k})^{1/\alpha}}\frac{(\log(\max I_{k}))^{\beta}}{(\log(\min I_{k}))^{\beta}}=\lim_{k\to\infty}\left(\frac{\min I_{k}}{\max I_{k}}\right)^{1/\alpha}\left(1+\frac{\log\frac{\max I_{k}}{\min I_{k}}}{\log\min I_{k}}\right)^{\beta}=0,$
and hence
$\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(S\cap[1,\max I_{k}])}=1-\lim_{k\to\infty}\frac{\#(S\cap[1,\min I_{k}])}{\#(S\cap[1,\max I_{k}])}=1.$
Using this convergence and (4.11) we obtain
$\begin{split}\overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A|S\right)\geq&\limsup_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap[1,\max I_{k}]}{\#(S\cap[1,\max I_{k}])}\\ \geq&\limsup_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}}{\#(S\cap[1,\max I_{k}])}\\ =&\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(S\cap[1,\max I_{k}])}\\ &\times\lim_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}}{\#(S\cap I_{k})}\\ =&\overline{d}(A|S).\end{split}$
The reverse inequality $\overline{d}(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A|S)\leq\overline{d}(A|S)$ is obvious. We have verified (1.7). Combining (1.6) and (1.7) yields (1.8). The proof of Theorem 1.4 is complete.∎
4.2. Proof of Theorem 1.1(b)
In principle we proceed along the line of §$\S$4.1, but there are a number of technical differences. Let $S\subset\mathbb{N}$ satisfy $\overline{d}(S)=0$ and $d_{\rm B}(S)>0$. We set $M_{0}=N_{0}=0$ for convenience, and take two sequences $(M_{k})_{k=1}^{\infty}$, $(N_{k})_{k=1}^{\infty}$ of positive integers such that $(N_{k})_{k=1}^{\infty}$ is strictly increasing and
$\lim_{k\to\infty}\frac{1}{N_{k}}\#({S\cap[M_{k},M_{k}+N_{k}]})=d_{\rm B}(S).$ (4.13)
If $\sup_{k\geq 1}M_{k}<\infty$ then we have $d_{\rm B}(S)=\overline{d}(S)$. This case is covered by Theorem 1.1(a). In what follows we assume $\sup_{k\geq 1}M_{k}=\infty$. Taking subsequences of $(M_{k})_{k=1}^{\infty}$ and $(N_{k})_{k=1}^{\infty}$ if necessary, we may assume $(M_{k})_{k=1}^{\infty}$ is strictly increasing and $M_{k}+N_{k}<M_{k+1}$ for every $k\geq 1$. For each $k\geq 1$, set
$I_{k}=[M_{k-1},M_{k-1}+N_{k-1}].$
We have $\lim_{k\to\infty}\#(\mathbb{N}\cap I_{k})=\infty$ and
$\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(\mathbb{N}\cap I_{k})}=d_{\rm B}(S).$ (4.14)
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a monotone decreasing sequence of positive reals converging to 00. Taking subsequences of $(M_{k})_{k=1}^{\infty}$ and $(N_{k})_{k=1}^{\infty}$ if necessary, we may assume
$\frac{1}{(M_{k-1}+N_{k-1}+1)^{2N_{k-1}}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1,$ (4.15)
and
$\inf_{n\geq M_{k}}\prod_{i=1}^{n+1}\frac{1}{(t^{i+2})^{2}}\cdot\left(\prod_{i=1}^{n}\frac{1}{t^{2i}}\right)^{-1-\varepsilon_{k}}\geq 2\ \text{ for every }k\geq 1.$ (4.16)
Then (4.15) implies
$\prod_{i\in S\cap I_{k}}\frac{1}{(i+1)^{2}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1\text{ with }S\cap I_{k}\neq\emptyset.$ (4.17)
Since $\overline{d}(S)=0$, the set $\mathbb{N}\setminus S$ has polynomial density with $\alpha=1$, $\beta=0$. By Proposition 3.1, there exist a subset $(\mathbb{N}\setminus S)_{*}$ of $\mathbb{N}\setminus S$ not containing consecutive integers and positive reals $t,L$ with $t>L>1$ and $t>2$ such that $R_{t,L}((\mathbb{N}\setminus S)_{*})$ is an extreme seed set of Hausdorff dimension $1/2$. We set
$F_{S}=R_{t,L}((\mathbb{N}\setminus S)_{*},S,(M_{k})_{k=1}^{\infty},(S\cap I_{k})_{k=1}^{\infty}).$
The construction of this set was described in §$\S$3.2. For each $y\in R_{t,L}((\mathbb{N}\setminus S)_{*})$ we add the elements of $S\cap I_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ at position $M_{k}$ to define a new sequence. Then $F_{S}$ is the set of points whose regular continued fraction expansions are given by these new sequences.
Since the intervals $I_{k}$, $k\geq 1$ are pairwise disjoint and $(\mathbb{N}\setminus S)_{*}\cap S=\emptyset$, we have $F_{S}\subset E$. Since (3.10) and (3.11) follow from (4.17) and (4.16) respectively, the corresponding elimination map is almost Lipschitz by Proposition 3.5. Lemma 2.4 gives
$\dim_{\rm H}F_{S}\geq\dim_{\rm H}R_{t,L}((\mathbb{N}\setminus S)_{*})=\frac{1}{2}.$ (4.18)
By the construction of $F_{S}$, we have
$\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\cap I_{k}=S\cap I_{k}.$
Then (4.14) yields
$\begin{split}d_{\rm B}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\right)&\geq\limsup_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\cap I_{k}}{\#(\mathbb{N}\cap I_{k})}\\ &=\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(\mathbb{N}\cap I_{k})}=d_{\rm B}(S).\end{split}$
The reverse inequality $d_{\rm B}(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S)\leq d_{\rm B}(S)$ is obvious. We have verified (1.4), which implies (1.5). The proof of Theorem 1.1(b) is complete. ∎
4.3. Proof of Theorem 1.13
Let $S$ be an infinite subset of $\mathbb{N}$. By Lemma 2.7, if $x\in E$ then $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0$ implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=1$ which in turn implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})=1$. Applying Lemma 2.8 with $c=1$ we conclude that $\left\{x\in J\colon\ \overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0\right\}$ is of Hausdorff dimension zero. So, if $\overline{d}(S)>0$ then $\left\{x\in J\colon\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)\right\}$ is of Hausdorff dimension zero. Since $\dim_{\rm H}J=1/2>0$ by Lemma 3.4, the desired inequality in Theorem 1.13(a) holds.
Suppose $S$ has polynomial density with exponent $\alpha<2$. Lemma 2.7 gives $\tau(S)=1/\alpha$. Let $A\subset S$ satisfy $\overline{d}(A|S)>0$. By Lemma 2.7, if $x\in E$ then $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)>0$ implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A)=1/\alpha$ which in turn implies $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})\geq 1/\alpha$. If moreover $\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S$ then $\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})\leq\tau(S)=1/\alpha$. Further, applying Lemma 2.8 with $c=1/\alpha$ we obtain
$\dim_{\rm H}\left\{x\in J\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S,\ \overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)>0\right\}\leq\frac{\alpha-1}{2\alpha}.$
Combining this inequality, Corollary 3.4 and then using $\alpha<2$ yields the desired inequality in Theorem 1.13(b).∎
Appendix A Bergelson-Leibman’s theorem in $\mathbb{N}$
This appendix gives the proof of Theorem 1.2. It relies on Furstenberg’s correspondence principle and the following multiple recurrence theorem with polynomial times. Recall that $\mathbb{Z}_{0}[X]$ denotes the set of polynomials with integer coefficients that vanish at $X=0$.
Theorem A.1 ([2] Theorem ${\rm A}_{0}$).
Let $(X,\mathcal{B},\mu)$ be a probability space, let $T_{1},\ldots,T_{\ell}$ be a finite collection of commuting invertible measure preserving transformations on $X$, let $h_{1},\ldots,h_{\ell}\in\mathbb{Z}_{0}[X]$, and let $A\in\mathcal{B}$ satisfy $\mu(A)>0$. Then
$\liminf_{N\to\infty}\frac{1}{N}\sum_{m=0}^{N-1}\mu(T_{1}^{-h_{1}(m)}(A)\cap T_{2}^{-h_{2}(m)}(A)\cap\cdots\cap T_{\ell}^{-h_{\ell}(m)}(A))>0.$
A.1. Proof of Theorem 1.2
Let $S\subset\mathbb{N}$ satisfy $d_{\rm B}(S)>0$. Let $\{0,1\}^{\mathbb{Z}}$ denote the two-sided Cartesian product space endowed with the product topology of the discrete topology on $\{0,1\}$. Define $x^{*}=(x_{i}^{*})_{i=-\infty}^{\infty}\in\{0,1\}^{\mathbb{Z}}$ by
$x_{i}^{*}=\begin{cases}1&\text{if $i\geq 1$ and $i\in S$,}\\ 0&\text{otherwise.}\end{cases}$
Let $\sigma$ denote the left shift acting on $\{0,1\}^{\mathbb{Z}}$ and let $\Sigma=\overline{\{\sigma^{i}(x^{*})\colon i\in\mathbb{Z}\}}$. Then $\Sigma$ is a compact metrizable space and $\sigma(\Sigma)=\Sigma$. Let $A=\{(y_{i})_{i=-\infty}^{\infty}\in\Sigma\colon y_{0}=1\}$. Let $(I_{k})_{k=1}^{\infty}$ be a sequence of compact intervals in $[1,\infty)$ such that #$\#(\mathbb{N}\cap I_{k})\to\infty$ as $k\to\infty$ and
$d_{\rm B}(S)=\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(\mathbb{N}\cap I_{k})}.$
For each $k\geq 1$ define a Borel probability measure $\mu_{k}$ on $\Sigma$ by
$\mu_{k}=\frac{1}{\#(\mathbb{N}\cap I_{k})}\sum_{i\in\mathbb{N}\cap I_{k}}\delta_{\sigma^{i}(x^{*})},$
where $\delta_{\sigma^{i}(x^{*})}$ denotes the unit point mass at $\sigma^{i}(x^{*})$. Note that $\lim_{k\to\infty}\mu_{k}(A)=d_{\rm B}(S)$. Pick a limit point of the sequence $(\mu_{k})_{k=1}^{\infty}$ in the weak* topology and denote it by $\mu$. Then $\mu$ is $\sigma|_{\Sigma}$-invariant. Since $A$ is a clopen set, we have $\mu(A)=\lim_{k\to\infty}\mu_{k}(A)$ and so $\mu(A)=d_{\rm B}(S)>0$.
For each $m\in\mathbb{Z}$ let $S-m=\{i\in\mathbb{N}\colon i+m\in S\}.$ Let $k\geq 1$ and $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$. For each $i\in\mathbb{N}$, note that $i\in I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell})$ if and only if $\sigma^{i}(x^{*})\in A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)$. This correspondence yields
#$\begin{split}\frac{\#(I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))}{\#(\mathbb{N}\cap I_{k})}=\mu_{k}(A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)).\end{split}$
Hence
$\begin{split}d_{\rm B}(S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))&\geq\lim_{k\to\infty}\frac{\#(I_{k}\cap S\cap(S-m_{1})\cap\cdots\cap(S-m_{\ell}))}{\#(\mathbb{N}\cap I_{k})}\\ &=\lim_{k\to\infty}\mu_{k}(A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A))\\ &=\mu(A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)).\end{split}$
The first inequality follows from the definition of upper Banach density. The second equality holds because $A\cap\sigma^{-m_{1}}(A)\cap\cdots\cap\sigma^{-m_{\ell}}(A)$ is a clopen set. Let $h_{1},\ldots,h_{\ell}\in\mathbb{Z}_{0}[X]$. Replacing $m_{j}$ by $h_{j}(m)$, $m\in\mathbb{N}$ for $j=1,\ldots,\ell$ we have
$d_{\rm B}(S\cap(S-h_{1}(m))\cap\cdots\cap(S-h_{\ell}(m)))\geq\mu(A\cap\sigma^{-h_{1}(m)}(A)\cap\cdots\cap\sigma^{-h_{\ell}(m)}(A)).$
Applying Theorem A.1 to the invertible $\mu$-preserving system $(\Sigma,\sigma|_{\Sigma})$ we obtain $d_{\rm B}(S\cap(S-h_{1}(m))\cap\cdots\cap(S-h_{\ell}(m)))>0$ for infinitely many $m\in\mathbb{N}$. This yields the desired statement.∎
Appendix B Arithmetic progressions in ${\rm PS}(\alpha)$
This appendix gives the proof of Theorem 1.10 for the sets obtained by the Piatetski-Shapiro sequences. Recall that ${\rm PS}(\alpha)=\{\lfloor n^{\alpha}\rfloor\colon n\in\mathbb{N}\}$ for $\alpha>1$.
B.1. Proof of Theorem 1.10
Let $21<\alpha<$ 2. Let $A\subset{\rm PS}(\alpha)$ satisfy $\overline{d}(A|{\rm PS}(\alpha))>0$. There exists a strictly increasing sequence $(N_{k})_{k=1}^{\infty}$ of positive integers such that for every $k\geq 1$,
#$\frac{\#(A\cap[1,N_{k}])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}\geq\frac{\overline{d}(A|{\rm PS}(\alpha))}{2}.$
Set $B=\{n\in\mathbb{N}\colon\lfloor n^{\alpha}\rfloor\in A\}$. For every $k\geq 1$ we have
#$\frac{\#(B\cap[1,\#({\rm PS}(\alpha)\cap[1,N_{k}])])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}=\frac{\#(A\cap[1,N_{k}])}{\#({\rm PS}(\alpha)\cap[1,N_{k}])}\geq\frac{\overline{d}(A|{\rm PS}(\alpha))}{2}>0.$
This implies $d_{\rm B}(B)\geq\overline{d}(B)>0$. By [32, Corollary 5], $\{(n,\lfloor n^{\alpha}\rfloor)\in\mathbb{N}^{2}\colon n\in B\}$ contains an $\ell$-term arithmetic progression in $\mathbb{N}^{2}$ for every $\ell\geq 3$. Therefore, $A=\{\lfloor n^{\alpha}\rfloor\colon n\in B\}$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.∎
Appendix C Generalizations to Iterated Function Systems
A countable collection of self-maps is called an Iterated Function System (IFS). It is well-known that the regular continued fraction (1.1) is generated by an IFS on $[0,1]$. Theorems 1.1 and 1.4 can be generalized to a wide class of expansions of real numbers in $[0,1]$ that are generated by IFSs.
Let $\{f_{k}\colon k\in\mathbb{N}\}$ be a collection of $C^{1}$ maps from $[0,1]$ to itself and let $d>1$. We say $\{f_{k}\colon k\in\mathbb{N}\}$ is a $d$-decaying IFS [18] if the following conditions hold:
  • (i)
    there exists $m\in\mathbb{N}$ and $\gamma\in(0,1)$ such that for every $(a_{1},\ldots,a_{m})\in\mathbb{N}^{m}$ and every $x\in[0,1]$ we have $0<|(f_{a_{1}}\circ\cdots\circ f_{a_{m}})^{\prime}(x)|\leq\gamma;$
  • (ii)
    for all distinct $i,j\in\mathbb{N},$ $f_{i}((0,1))\cap f_{j}((0,1))=\emptyset;$
  • (iii)
    for any $\varepsilon>0,$ there exist $c_{1}(\varepsilon)$, $c_{2}(\varepsilon)$ with $0<c_{1}(\varepsilon)\leq 1\leq c_{2}(\varepsilon)$ such that for every $n\in\mathbb{N}$ and every $x\in[0,1]$ we have
    $\frac{c_{1}(\varepsilon)}{k^{d(1+\varepsilon)}}\leq|f_{k}^{\prime}(x)|\leq\frac{c_{2}(\varepsilon)}{k^{d(1-\varepsilon)}};$
  • (iv)
    $\overline{\bigcup_{k=1}^{\infty}f_{k}([0,1])}=[0,1]$, and if $i<j$ then $f_{i}(x)>f_{j}(x)$ for all $x\in[0,1].$
If $\{f_{k}\colon k\in\mathbb{N}\}$ is a $d$-decaying IFS, then by (i) one can define a natural projection $\Pi\colon\mathbb{N}^{\mathbb{N}}\rightarrow[0,1]$ by
$\Pi((a_{n})_{n=1}^{\infty})=\lim_{n\to\infty}f_{a_{1}}\circ\cdots\circ f_{a_{n}}(1).$
For any $x\in[0,1],$ there exists a sequence $(a_{n})_{n=1}^{\infty}$ of positive integers such that
$x=\Pi((a_{n})_{n=1}^{\infty}).$ (C.1)
By (ii), this correspondence is one-to-one up to a countable number of points, and (C.1) gives an expansion of $x$ into an infinite series. The sequence $(a_{n})_{n=1}^{\infty}$ in (C.1) may be viewed as a digit sequence of $x$.
The regular continued fraction is generated by the $2$-decaying IFS $\{f_{k}\colon k\in\mathbb{N}\}$ defined by $f_{k}(x)=1/(x+k)$ with $m=2$. Our setup and main results can be generalized to number expansions generated by arbitrary $d$-decaying IFSs. The constant $1/2$ in Theorem 1.1 should be replaced by $1/d$, and the constant $1/(2\alpha)$ in Theorem 1.4 should be replaced by $1/(d\alpha)$. Lemma 2.1 is the only place where the specific formulas for the regular continued fraction were used. In this generalization, (iii) can be used as a substitute for Lemma 2.1.
Let us stick to Billingsley’s notation for continued fractions. Each $x\in(0,1)\setminus\mathbb{Q}$ has the unique continued fraction expansion of the form
[…]
where $a_{n}\in\mathbb{N}$ for $n\in\mathbb{N}$, which is called the minus or backward continued fraction [16]. This continued fraction is generated by the IFS $\{f_{k}\colon k\in\mathbb{N}\}$ on $[0,1]$ defined by $f_{k}(x)=1-1/(x+k)$ that does not satisfy (i) since $f_{1}(0)=0$ and $f_{1}^{\prime}(0)=1$. Our main results can be generalized to this expansion and the proofs remain intact, since only those $x$ with $a_{n}\to\infty$ as $n\to\infty$ is concerned in the setup. Even further, our main results can be generalized to semi-regular continued fractions that include the regular and the backward continued ones as special cases, see [26] for details.
Acknowledgments
We thank Yasushi Nagai and Kota Saito for their careful readings of an earlier version of the draft and giving valuable suggestions. We thank Nima Alibabaei, Atsushi Katsuda, Shunsuke Usuki for fruitful discussions. This research was supported by the JSPS KAKENHI 23K20220.